a.A source speech signal is encoded by an 8th-order Linear Predictive Coding (LPC) parametric model using...

  • a.A source speech signal is encoded by an 8th-order Linear Predictive Coding (LPC) parametric model using a frame analysis length of 20ms and with a 10% frame overlap.

    How many frames per second will be encoded by the LPC model?

    (2 marks)

  • b.Following quantising in the LPC model described in part a., an average of 5 bits per frame is allocated to transmit each predictor coefficient. If it is assumed that 1 bit is allocated for the input excitation (U/V) signal, and 3 bits each are allocated to both the pitch period and amplitude gain control, what will be the output bit-rate of the LPC model?

    (3 marks)
Answer #1

a. 10% of 20 ms = 20 *0.1 ms = 2 ms

i.e. Two consecutive frames have 2 ms overlap

No of frames in 1 sec

= floor(1000 ms /(20 - 2 )ms)

= floor(55.5556)

= 55 [Ans]

b. 8th order LPC has 8+1=9 predictor coefficients per frame.

Predictor coefficients generated from each frame are represented by 9 x 5 = 45 bits

Adding the bits for input excitation signal, pitch period and amplitude gain control, total number of bits to represent each frame

= 45 + 1 + 3 + 3

= 52

Bit rate

= 52 bits/frame x 55 frames/second

= 2860 bits/second